OpenStudy (anonymous):

Solution A is 10% sugar while solution b is 30% sugar. A chemist have 60-liter mixture of solutions A and B with 18% sugar. how many liters of each solution should he use?

OpenStudy (anonymous):

I need the formula also plsss

OpenStudy (anonymous):

sorry to interput but is anyone good with inverse functions? :?

OpenStudy (trancenova):

C1V1=C2V2 Where '2' is the final solution C2V2 = C1V1 + C3V3 (we don't know V1 and V3) 0.18x60=0.3xV1+0.1xV3 V1+V3=60 V1=60-V3 0.18x60=0.3(60-V3)+0.1V3 0.18x60=-0.3V3 +18 + 0.1V3 0.18x60=-0.2V3 +18 0.18x60-18 = -0.2V3 10.8 + 18 =-0.2V3 -7.2 = -0.2V3 7.2 = 0.2V3 7.2/0.2 =V3 V3= 7.2/0.2 =36 10% solution (A) needs 36L in the 60L to make a 18% soln. So you need 36L of A and 24L of B.

OpenStudy (anonymous):

thanks for the help TraceNova Group Lifesaver